if A-B is extended its own length at what point would it extend. A (-4,4) to B (2,8). what is the formula. the professor said something about taking the averages of the x y values.
Answer by Mathtut(3670)

yes the midpoint formula states that between any two points the midpoint is
:
((x1+x2)/2,(y1+y2)/2).....so the average of the x coordinates and the average of the y coordinates
:
In this case we know what the midpoint is(point B). What we are after is the other endpoint lets call it C=(x,y)
:
so we have ((-4+x)/2=2 , (4+y)/2=8)
:
-4+x=4
:
x=8
:
4+y=16
:
y=12
:
so it would extend to C= (8,12)